OpenStudy (anonymous):

0.5 as a decimal and fraction in lowest form 0.5 as a decimal and fraction in lowest form @Mathematics

OpenStudy (chriss):

1/2

OpenStudy (anonymous):

and 0.50

OpenStudy (saifoo.khan):

same.

OpenStudy (anonymous):

o.o but the decimAl

OpenStudy (saifoo.khan):

50/100 = 1/2 5/10 = 1/2

OpenStudy (anonymous):

O.o thanks... And the decimal is 0.50 right

OpenStudy (saifoo.khan):

hmm

OpenStudy (zarkon):

\[.4\overline{9}\]